OpenStudy (anonymous):

Derive this identity from the sum and difference formulas for cosine: sin a sin b = (1 / 2)[cos(a – b) – cos(a + b)]

OpenStudy (anonymous):

Start with the right hand side because it is more complex

OpenStudy (anonymous):



OpenStudy (anonymous):

@tkhunny please help

OpenStudy (tkhunny):

Please write down the sum and difference formulas for cosine.

OpenStudy (anonymous):

Cos(a + b) = cos a cos b + sin a sin b sin (a + b) = sin a cos b + sin b cos a

OpenStudy (tkhunny):

No, that's the sum formula for sine and cosine. Write down the sum and difference formulas for cosine. We don't want any such thing for sine. Actually, that's not the sum formula for cosine, either. Look at the cosine formulas. Look at them very carefully. Get them right.

OpenStudy (anonymous):

cos(a-b) = cosA cosB + sinA sinB cos(a+b) = cosA cosB - sinA sinB ?

OpenStudy (tkhunny):

That's better. Question: What will happen if we just add these two equations?

OpenStudy (anonymous):

You'd get 0. Don't they cancel each other out?

OpenStudy (tkhunny):

No. You're guessing. You need to start proving. If we add them, we get... cos(a-b) + cos(a+b) = 2cos(A)cos(b) The parts with the sine do go away. Reassess. How close are we? Are we going the right way? If not, can we fix it?

OpenStudy (anonymous):

Well it looks like we are getting closer. We have cos(a-b) + cos(a+b) which is similar to the cos(a-b) - cos(a+b)

OpenStudy (tkhunny):

It's a good try, but we need the sine portions, not the cosine portions. What happens if we SUBTRACT, rather than add?

OpenStudy (anonymous):

Truthfully I'm not sure. I know the beginning will be cos(a-b) - cos (a+b) = …

OpenStudy (tkhunny):

Right, then the cosine pieces go away. What happens with the sine pieces?

OpenStudy (anonymous):

2sin(a)sin(b)

OpenStudy (anonymous):

or -2

OpenStudy (tkhunny):

cos(a-b) = cos(a) cos(b) + sin(a) sin(b) cos(a+b) = cos(a) cos(b) - sin(a) sin(b) Subtracting cos(a-b) - cos(a+b) = 2sin(a) sin(b) Reassess. How close are we? Are we going the right way? If not, can we fix it?

OpenStudy (anonymous):

I believe we are because now we have the cos(a-b) - cos(a+b) which is needed

OpenStudy (anonymous):

@niksva pls help

OpenStudy (anonymous):

@washcaps, @tkhunny has explained this problem really well

OpenStudy (anonymous):

but I still don't know how to fill in the table which is the question

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

we are using the formulae for cos(A+B) and cos(A-B), so that can be put in the reasons group as far as the calculation is concerned , we have just done the subtraction i dont think we need to give the proof of cos(A+B) or other one

OpenStudy (anonymous):

so. 1. I will show what I will subtract ---> reason use formula cos(a+b) and cos(a-b)?

OpenStudy (anonymous):

That doesn't sound like the right reason @niksva

OpenStudy (anonymous):

this is mathematics we use formulae if u want to give more reasons , then give them the proof of cos(A+b) and cos(A-B)

OpenStudy (anonymous):

This would be easier if you could lay it out all in a table. I'm getting quite confused. I would really appreciate it @niksva

OpenStudy (anonymous):

Sorry,i could have but it is not allowed here (as per the rules of this forum) people are here to help only u have to do your work

OpenStudy (anonymous):

Well could you explain a little better. I'm not asking for you to hand me the answers. I want to see how you actually got the answers. I know the first calculation is the subtracting, but I don't know the reason or what comes after @niksva

OpenStudy (anonymous):

if i had known the better way, i would have explained it to you @tkhunny has explained all the steps to you really well there is nothing which I can add to his explanation

OpenStudy (anonymous):

I only see why step that @tkhunny explained. There are supposed to be 3

OpenStudy (anonymous):

not why. I meant one

OpenStudy (anonymous):

i will write the steps, cos(a-b) = cosa*cosb+sina*sinb ------(1) cos(a+b) =cosa*cosb - sina*sinb -------(2) (1) - (2) cos(a-b) - cos(a+b) = 2sina*sinb divide both sides by 2 (1/2)*{cos(a-b)-cos(a+b)} = sina*sinb

OpenStudy (anonymous):

so 1. cos(a-b) = cosa*cosb+sina*sinb -----> reason: use cosine sums and differences cos(a+b) =cosa*cosb - sina*sinb 2. cos(a-b) - cos(a+b) = 2sina*sinb ----> reason: subtracted difference 3. (1/2)*{cos(a-b)-cos(a+b)} = sina*sinb ----> reason: ???? @niksva

OpenStudy (anonymous):

divide both sides by 2

OpenStudy (anonymous):

Does this seem like a correct answer to fill in the table with?

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

Thank you! I hope it is right lol